Project

General

Profile

Bug #1184

Repair status

Added by tin over 4 years ago. Updated over 3 years ago.

Status:
Closed
Priority:
Normal
Assignee:
Target version:
Start date:
11/04/2015
Due date:
% Done:

100%

Estimated time:

Description

Analog board A14764
Digital board A39541 B15

Fixed eaten trace under caps.
Replaced all caps

Self-test issues:

Test A/D input Required value
401.1 -76m +/-1m
402.1 +25.0m 27 - 37 m
404.1 5.5705 3.81 - 4.61
404.2 4.1110 1.74 - 2.42
404.3 2.7505 279 - .281
404.4 1.3644 1.91 - 2.59
404.5 0.1282 3.93 - 4.73
405.2 11.8598
405.4 11.8592
405.6 11.8588
405.8 11.8584
406.6 11.8610
407.1 2.2620 6.85 - 7.15
407.2 11.8592 0.5 - 0.9
407.3 8.9122 0.1 - 0.18
409.6 6.8057
410.1 2.2621 6.82 - 7.18
411.1 2.2623 6.82 - 7.18
411.2 2.2629 6.84 - 7.16

History

#1

Updated by tin almost 4 years ago

Diagnostics revealed function issues - DCI measurement on ranges other than 2A was throwing some incorrect readings even with short on inputs, ACV was showing 480VAC on 750V range (no input signal) and overflow on lower ranges and ACI currents. Few other DG211 muxes, 14094 shift registers were replaced during debug process, but that was likely not necessary.

Issues fixed after U525 and 4.3V zener replacement. Original LTC1050 opamp had leaky 145 ohm between V+ and V- rails making floating supply circuitry unbalanced.

Second meter have bad LT1223 current mode operational amplifier, driving FET full-wave rectifier bridge to -12V, and screwing everything. Same dianosis errors and ACV,DCI and ACI function failures were observed. Removed U520 and powered meter without it, supply voltages on U525 become normal.

#2

Updated by tin over 3 years ago

  • Status changed from In Progress to Resolved
  • % Done changed from 0 to 100
#3

Updated by tin over 3 years ago

  • Status changed from Resolved to Closed

Also available in: Atom PDF